OpenStudy (aravindg):

If A+B+C=180 prove that cos^2a+cos^2B+cos^2C=1-COSACOSBCOSC

OpenStudy (aravindg):

SRY 1-2COSACOSBCOSC

OpenStudy (aravindg):

helppppppp

OpenStudy (anonymous):

If A+B+C=180, then A+B=180-C, and hence \[\cos(A+B) = \cos(180-C) = -\cos C\] Also, we're going to use the following trigonometric identities: \begin{eqnarray*} \cos x\cos y &=& \cos(x+y) + \cos(x-y) \\ \cos 2x &=& 2\cos^2x-1 \end{eqnarray*} Let's start from 1 - 2cosAcosBcosC. \begin{eqnarray*} 1-2\cos A\cos B\cos C&=&1-\cos C(\cos A\cos B) \\ &=&1-\cos C(\cos(A+B) + \cos(A-B)) \\ &=&1-\cos C(-\cos C + \cos(A-B)) \\ &=&1+\cos^2C - \cos C\cos(A-B) \\ &=&1+\cos^2C + \cos(A+B)\cos(A-B) \\ &=&1+\cos^2C + \frac{1}{2}(\cos(A+B+A-B)+\cos(A+B-A+B)) \\ &=&1+\cos^2C + \frac{1}{2}(\cos 2A + \cos 2B) \\ &=&1+\cos^2C + \frac{1}{2}(2\cos^2A-1+2\cos^2B-1) \\ &=&1+\cos^2C + \frac{1}{2}(2\cos^2A+2\cos^2B-2) \\ &=&1+\cos^2C + \cos^2A + \cos^2B - 1 \\ &=&\cos^2A+\cos^2B+\cos^2C \end{eqnarray*} as required.

OpenStudy (aravindg):

wow another one

OpenStudy (aravindg):

If A+B+C=180 prove that tan (A/2)*tan(B/2)+tan(B/2)*tan(C/2)+tan(C/2)*tan(A/2)=1

OpenStudy (aravindg):

xactxx one doubt 2cosxcosy=cos(X+y)+cos(x-y) and not cosx cosy=cos(X+y)+cos(x-y)

OpenStudy (anonymous):

You're right, there should be a 2 on the LHS or 1/2 on the RHS

myininaya (myininaya):

xactxx fantastic job!

OpenStudy (aravindg):

plz rectify it

OpenStudy (anonymous):

How do I edit my own post?

OpenStudy (aravindg):

myininaya y u say fantastic ??????myininaya he made a mistake

myininaya (myininaya):

but still it is an awesome job

OpenStudy (aravindg):

gn8 guys ..plz post the answer

myininaya (myininaya):

i wouldn't have thought to do A+B=180-C then say cos(A+B)=cos(180-C)=-1cos(-C)=-cos(-C)=-cos(C)

OpenStudy (anonymous):

is there a way to edit my post please?

myininaya (myininaya):

no

myininaya (myininaya):

but you can copy the latex above and post it again and change something if you want

myininaya (myininaya):

right click and then say show source

OpenStudy (anonymous):

Well I'll just repost the answer with the fixes. If A+B+C=180, then A+B=180-C, and hence \[\cos(A+B)=\cos(180−C)=−\cos C\] Also, we're going to use the following trigonometric identities: \begin{eqnarray*} 2\cos x\cos y &=& \cos(x+y)+\cos(x-y) \\ \cos 2x&=&2cos^2x−1 \end{eqnarray*} Let's start from 1 - 2cosAcosBcosC. \begin{eqnarray*} 1−2\cos A\cos B\cos C&=&1−\cos C(2\cos A\cos B) \\ &=&1−\cos C(\cos(A+B)+\cos(A−B)) \\ &=&1−\cos C(−\cos C+\cos(A−B)) \\ &=&1+\cos^2C−\cos C\cos(A−B) \\ &=&1+\cos^2C+\cos(A+B)\cos(A−B) \\ &=&1+\cos^2C+\frac{1}{2}(\cos(A+B+A−B)+\cos(A+B-A+B)) \\ &=&1+\cos^2C+\frac{1}{2}(\cos 2A+\cos 2B) \\ &=&1+\cos^2C+\frac{1}{2}(2\cos^2A−1+2\cos^2B−1) \\ &=&1+\cos^2C+\frac{1}{2}(2\cos^2A+2\cos^2B−2) \\ &=&1+\cos^2C+\cos^2A+\cos^2B−1 \\ &=&\cos^2A+\cos^2B+\cos^2C \end{eqnarray*} as required.